1)(a) Find the limit of the sequence whose kth term is sirk) Confirm your answer graphically, and show your graph for full cred

2)Use formulas for finite geometric series and infinite geometric series to find exact answers for each of the following: т S 6

3)Use vertical rectangle slices to write a definite integral which is equal to the area of the triangular region with comers at

4)In some situations, the force required to stretch a “spring is not linear. Consider the following model of a bow and arrow i

5)Drugs that decay exponentially leave a residue in the body. By definition, the half- life of a drug is the time until half th
Answer #1

Solution:a)To find the limit of the sequence whose k^{th} term is sin(k) , we need to solve the following:

\lim_{k\rightarrow \infty}\frac{\sin (k)}{k}

The value of \sin (k) alternates between -1 and 1 irrespective of the value of  k. Therefore, the numerator term i.e.  \sin (k) of the given series will always have a finite value, whereas the value of denominator term keeps on increasing and finally tends to infinity. Therefore, when the value of k is substituted in the given limit, we will get 0 as result. Therefore, we get:

\boxed{\lim_{k\rightarrow \infty}\frac{\sin (k)}{k}=0}

In order to verify this graphically, graph the function  sin(k) using any graphical tool. We will get the following graph:

As can be seen from the graph above, the value of the given sequence keeps on decreasing progressively as the value of k increases and will ultimately die down to zero.

b) Given to us that the sequence b_k converges:

b_k=\frac{k^2+4k}{3k^2+7k+8}

In order to find the limit of the given convergent series, we will divide the numerator and the denominator terms by k^2 to get:

\lim_{k\rightarrow \infty}b_k=\lim_{k\rightarrow \infty} \ \frac{1+\frac{4}{k}}{3+\frac{7}{k}+\frac{8}{k^2}}

Substituting the value of limit, we get:

=\frac{1+\frac{4}{\infty}}{3+\frac{7}{\infty}+\frac{8}{\infty}}

=\frac{1+0}{3+0+0}

\Rightarrow \boxed{\lim_{k\rightarrow \infty}=\frac{1}{3}}